# MagicFace: High-Fidelity Facial Expression Editing with Action-Unit Control

### v2

@@ Abstract
We address the problem of facial expression editing by controling the relative variation of facial action-unit (AU) from the same person. This enables us to edit this specific person’s expression in a fine-grained, continuous and interpretable manner, while preserving their identity, pose, background and detailed facial attributes. Key to our model, which we dub MagicFace, is a diffusion model conditioned on AU variations and an ID encoder to preserve facial details of high consistency. Specifically, to preserve the facial details with the input identity, we leverage the power of pretrained Stable-Diffusion models and design an ID encoder to merge appearance features through self-attention. To keep background and pose consistency, we introduce an efficient Attribute Controller by explicitly informing the model of current background and pose of the target. By injecting AU variations into a denoising UNet, our model can animate arbitrary identities with various AU combinations, yielding superior results in high-fidelity expression editing compared to other facial expression editing works. Code is publicly available at https://github.com/weimengting/MagicFace.
Index Terms:
Facial Action unit, Facial expression editing, Diffusion models.
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
It is a perennial challenge in computer vision to realistically change the expression of a closeup while preserving the person’s identity and other attributes either from background or other face characteristics. The challenge of this problem arises from the lack of intuitive and interpretable depiction to represent facial expressions that can support customized expressions, and previous work usually addresses this with a latent space of expressions, where the codes are learned from a large expression dataset or the off-the-shelf ones like 3DMM parameters [1, 2, 3, 4, 5]. These methods ignore the fact that the semantic meanings of these codes are implicit, which poses challenges for interpretable, arbitrary and flexible manupulation of expression by non-professionals. In this work, we show that it’s possible to convincingly alter a person’s expression, in a user-friendly manner by offering localized control with adjustable intensity, while preserving their identity and other attributes from the portrait. Our key insight is to employ action units (AUs) to represent facial expressions, and then steer a Stable-Diffusion model to produce high-quality expression editing results.
@@
In editing facial expressions, the crucial question is which representation should be used to encode expressions. Facial AUs, as anatomical markers of facial muscle activity, have proven effective for precise and flexible facial expression description in images [6, 7, 8]. However, most AU annotations are of frontal faces of limited number of subjects in laboratory settings, which is easy to cause overfitting when used for training. A model trained by this type of data is hard to generalize on individuals with different poses and backgrounds, which we will present in Section IV-C. To deal with this issue, off-the-shelf AU intensity estimators like Libreface [9] used a knowledge distillation technique from a large-scale network pre-trained on natural images to accommodate AU intensity estimation task on the lab datasets. Such automatic tools provide us with AU intensity estimation of high accuracy, hence we can produce AU estimation of any face image and use the estimation as AU condition to train our model.
@@
On the other hand, diffusion models have emerged as a popular choice for image generation, surpassing Generative Adversarial Networks (GANs) with higher generation quality [10, 11, 12, 13]. ControlNet-style models enable users to add additional control signals such as depth, paintings, skeleton pose as the condition of the target [14]. Some works modify it to enable some basic expressions like happiness and sad [15, 14], but many of their editing results are with extreme face deformations that may look unrealistic. More importantly, these models can not provide specific controls over expressions, whether through text or through images. Such specific controls, including intensity and location of expressions, are the focus of our work.
@@
To merge the advantages of both worlds, we propose MagicFace, a model that allows to correlate AU changes to facial expressions and then produce a photorealistic edited image conditioned on AU variations. Specifically, MagicFace first extracts AU intensities from portrait photographs using an off-the-shelf method, perform desired AU changes, and finally uses a Stable-Diffusion model to map the AU variations into photorealistic images. As the edited images should preserve the background and pose, we first cut out the background of the portrait and draw the contour of the pose, which is learned by an Attribute Controller so that MagicFace only needs to perform conditional inpainting to the face. To maintain the consistency of identity and high-frequency facial characteristics, we introduce an ID encoder which is a symmetrical UNet structure to capture spatial details of the input indentity. This design allows the model to understand the relationship with the identity image within a uniform feature space, greatly enhancing the preservation of detailed appearance features. Our model is trained on a designed dataset with 30K image pairs. Fig. 1 shows the editing results of various pepole, background and AU combinations.
In summary, our contibutions are:
@@
  • An ID encoder capable of well preserving the attributes of the person in the edited portraits, unconstrained by any head poses, backgrounds and characters.
  • Quantitative and qualitative results show that our method presents a more interpretable, flexible and user-friendly editing manner with higher image generation quality compared to other facial expression editing methods.
@@ II Related Works
Figure 2: A display showcasing various action units and their corresponding intensity scales. Only a set of commonly used AUs are displayed here. For a complete collection of AUs with descriptions see [16].
@@ II-A Facial Action Units
The Facial Action Coding System (FACS) [17] is one of the most impactful methods for analyzing facial behavior. It is a comprehensive, anatomy-based system capable of encoding diverse facial movements through combinations of fundamental Action Units (AUs). AUs represent specific facial configurations resulting from the contraction of one or more facial muscles and are not influenced by emotional interpretation. The earlier version of FACS [17] included 44 Action Units (AUs), with 30 of them anatomically linked to specific facial muscles, while the remaining 14 were categorized as miscellaneous actions. In a later version [18], the criteria were updated: AU25, AU26, and AU27 were merged based on intensity, as were AU41, AU42, and AU43. The intensities of AUs can be assessed using a five-point ordinal scale, typically labeled with uppercase letters A to E, representing minimal to maximal intensity alongside the AU number. For example, AU2A indicates the minimum intensity of AU2, while AU2E represents its maximum intensity. However, this five-point scale is not uniform; for instance, levels C and D encompass a broader range of facial changes compared to levels A, B, and E. Fig. 2 shows example images labeled with specific AUs. The precision and clarity of AUs provide unmatched control over facial expressions, offering users an intuitive and easily interpretable set of tools, so we choose to use it as the condition to describe and edit facial expressions.
@@ II-B Facial Expression Editing
Facial expression editing is a challenging task as it demands a deep understanding of input facial images and prior knowledge about human expressions. Unlike general facial attribute editing, which primarily focuses on modifying the appearance of specific facial regions, facial expression editing is more complex as it often involves significant geometric transformations and simultaneous changes across multiple facial components. Remarkable progress has been made in earlier years with the advancement of GANs. ExprGAN [19] introduces an expression controller to adjust the intensity of generated expressions, but it relies on a pre-trained face recognizer to maintain identity information. StarGAN [20], on the other hand, enables image translation across domains using a single model and preserves identity features by minimizing a cycle loss. However, it is limited to generating only discrete expressions. GANimation [21] provides a more fine-grained method for editing facial expressions by leveraging Action Units (AUs). Similarly, ICface [22] utilizes AUs to enable controllable facial expressions in facial reenactment. Recent focus has shifted to diffusion models due to its ability to produce high-quality images. However, many text-to-image diffusion models are unable to support flexible facial expression editing and, in most cases, can only handle simple emotion labels. Instead, some approaches choose the expression coefficients of 3DMM as the condition to enable fine-grained control over facial expression. However, the large number of coefficients poses a significant challenge, as only experts can manually adjust the desired expressions, which limits its applicability for general user-level applications. In contrast, this work focuses on injecting AUs into pre-trained large diffusion models to enable more flexible and user-friendly editing of face.
@@ III Method
Figure 3: Overview of MagicFace. During training, a pair of images with the same identity but different pose, background as well as expressions is used respectively as the identity image and the target. AU variations are computed by an estimator and then sent into the denoising UNet as AU condition. Pose and background of the target is parsed into an image condition independently, dealed with an Attribute Controller and then inputted to the denoising UNet. ID encoder takes in the encoded identity image to edit for target AUs, where features in each transformer blocks are merged into the corresponding ones of the denoising UNet via self-attention. During inference, the conditional image will be parsed from the identity image.
@@ III-A Preliminariy
Stable Diffusion. Both the denoising UNet and ID encoder of our method inherit the same architecture as well as parameters of the denoising UNet in Stable Diffusion (SD). SD is developed from latent diffusion model (LDM) [23], which introduces learning of feature distributions to reduce computational complexity. It contains a fixed autoencoder to map input images into feature maps with smaller size. Formally, for an image $\mathbf{x}$ given, the encoder maps it into a latent representation via $\mathbf{z}=\mathcal{E}(\mathbf{x})$ and the decoder reconstructs it by $\hat{\mathbf{x}}=\mathcal{D}(\mathbf{z})$. The denoising UNet in SD learns to denoise noise $\boldsymbol{\epsilon}$ which is normally distributed into $\mathbf{z}$. During training, the latent map $\mathbf{z}$ is added Gaussian noise in $t$ timesteps and produce noised latent $\mathbf{z}_{t}$. The UNet is trained to predict the noise added, optimized by the following objective:
@@
where $\epsilon$ denotes the denoising UNet and $c$ represents condition embeddings. In original SD, $c$ is produced by a text encoder CLIP ViT-L/14 [24] to enable text-to-image generations. During inference, $\mathbf{z}_{T}$ is sampled from a normal distribution as the noised latent at timestep $T$, and then it is denoised iteratively into $\mathbf{z}_{0}$ by deterministic sampling process like DDPM [25], DDIM [26]. Then $\mathbf{z}_{0}$ will be reconstructed by the decoder into a generated image.
AU variations VS absolute AUs. Stepping back and reconsidering the conditioning inputs, using absolute AUs should be an intuitive choice. However, this condition configuration has a drawback: the model must estimate the actual AUs of the input image to decide whether to edit it. From an application standpoint, we are required to supply a value that must exactly match the corresponding AU in the source image even if no changes are intended. AU variations, as opposed to absolute AUs, represent the intended change in specific action units. This aligns with the definition of action units, which indicates the activation state of facial muscles. Therefore, in our model, we choose the difference between target AUs $\mathbf{c}_{tgt}$ and source AUs $\mathbf{c}_{ID}$ as the input condition, denoted as:
@@ III-B Architecture
Attribute Controller. As shown in Fig. 3, Attribute Controller takes in the latent representation of an image where the face is masked and the face pose is outlined by landmarks. This design is to solve the ill-posed problem: in videos shot of wild settings, there are no image pairs with identical backgrounds and poses, while expressions undergo significant changes for the same person, which means no ground truth of the image with target AUs is available. However, we can find numerous images of the same person with different backgound, pose and facial expressions. As an alternative, we separate the background and pose as an independent condition of the identity by parsing out the background image, which will avoid the model from learning all the information through the input identity image. On the parsed background image, we draw the contour to highlight the head pose in case that some backgrounds maybe black. It is theoretically possible that the denoising UNet will learn background and pose from the identity. Empirically, we find the network learns to use the Attribute Controller for background and pose information and does not rely on the ID encoder, we hypothsize this is because the conditional image is pixel-aligned with the ground truth and thus more easily used by the model.
@@
Attribute Controller is a convolution layer with 4 $\times$ 4 kernels, 2 $\times$ 2 strides, and 4 channels to align the latent of conditional image with the same resolution as the noise input. The processed conditional image is then appended to the noisy latent together as the input to the denoising UNet.
@@
ID encoder. We first evaluated the vanilla ControlNet [14] for identity preservation. As illustrated in Fig. 6, we observed that ControlNet struggles to preserve appearance consistency when generating human images coherent to AU prompt, making it unsuitable for our editing task. Many approaches employ the CLIP image encoder [27] to encode image conditions, but in our experiments it doesn’t adequately tackle problems associated with maintaining consistency in details. CLIP is designed to align semantic, high-level features with text, as a result it will lose many details in the generated images. On the other hand, recent studies [28, 29, 30] have highlighted the significant role of self-attention layers in diffusion models in maintaining the details of generated images. Building on these findings, we conducted an experiment on self-attention for identity control. In this setup, both the identity image and the noisy image are processed through the self-attention layers. A key observation from this architecture is that it inherently facilitates appearance resemblance between the two images. One possible explanation is that the self-attention layers in the UNet play a crucial role in spatially transmitting appearance information. As a result, they can function as a deformation module, enabling the generation of visually similar images with varying geometric structures.
@@
Given the above observations, we design a feature extraction encoder $\omega_{\theta}$ for the identity image which is aimed to preserve detailed face attributes, inspired by [29]. This encoder shares the same architecture with the denoising UNet $\phi_{\theta}$. Specifically, let $\mathbf{s}=\mathcal{E}(\boldsymbol{I}_{id})$ denote the encoded latent feature sent to the ID encoder $\omega_{\theta}$ where $\boldsymbol{I}_{id}$ is the identity image. The features extracted from $\omega_{\theta}$ will be merged into the denoising UNet $\epsilon_{\theta}$ through self-attention. Along with AU conditions, the ID encoder and the denoising UNet are jointly optimized by:
@@
where $\mathbf{g}=\mathcal{E}(\boldsymbol{I}_{bg})$ denotes the encoded latent feature sent to the Attribute Controller $\phi_{\theta}$, $\boldsymbol{I}_{bg}$ is the background image and $\mathbf{c}_{AU}$ represents the encoded AU condition. We omit the AU encoder in the framework as it’s simply a linear layer to map the input AU into a feature with the same dimension as the time conditions in SD (details are deferred to Sec. IV-C). The AU condition is added to the time-embedding condition in $\phi_{\theta}$. To merge identity features from $\omega_{\theta}$ into $\phi_{\theta}$ with self-attention, we first dive into the transformer block [31] of the ID encoder and denoising UNet, identifying each pair of feature maps that serve as inputs for the respective self-attention layers. For ease of explanation, we temporarily use $\mathbf{a}_{n}\in\mathbb{R}^{h_{n}\times w_{n}\times c_{n}}$ and $\mathbf{b}_{n}\in\mathbb{R}^{h_{n}\times w_{n}\times c_{n}}$ to denote the $n$th pair of feature maps from ID encoder and denoising UNet. In the denoising UNet, in each block the feature maps from both sides are first concatenated by $\mathbf{b}_{\mathbf{a}_{n}}=\mathbf{a}_{n}\bigoplus\mathbf{b}_{n}\in\mathbb{R}^{h_{n}\times 2w_{n}\times c_{n}}$, then we replace $\mathbf{b}_{n}$ with $\mathbf{b}_{\mathbf{a}_{n}}$ as the input to the self-attention layer. After coming out from self-attention layer, we crop out the fist half of the output feature map into $\mathbf{c}_{n}$ of resolution $h_{n}\times w_{n}\times c_{n}$ before input to the subsequent cross-attention layer.

### v3

@@ Abstract
We address the problem of facial expression editing by controlling the relative variation of facial action unit (AU) from the same person. This enables us to edit this specific person’s expression in a fine-grained, continuous, and interpretable manner, while preserving their identity, pose, background, and detailed facial attributes. Key to our model, which we dub MagicFace, is a diffusion model conditioned on AU variations and an ID encoder to preserve facial details of high consistency. Specifically, to preserve the facial details with the input identity, we leverage the power of pretrained Stable-Diffusion models and design an ID encoder to merge appearance features through self-attention. To keep background and pose consistency, we introduce an efficient Attribute Controller by explicitly informing the model of the current background and pose of the target. By injecting AU variations into a denoising UNet, our model can animate arbitrary identities with various AU combinations, yielding superior results in high-fidelity expression editing compared to other facial expression editing works. Code is publicly available at https://github.com/weimengting/MagicFace.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ I Introduction
It is a perennial challenge in computer vision to realistically change the expression of a close-up while preserving the person’s identity and other attributes, either from the background or other face characteristics. The challenge of this problem arises from the lack of an intuitive and interpretable depiction to represent facial expressions that can support customized expressions, and previous work usually addresses this with a latent space of expressions, where the codes are learned from a large expression dataset or the off-the-shelf ones like 3DMM parameters [yan2019joint, zhu2024stableswap, liang2019adaptive, hou2022textface, peng2011editing]. These methods ignore the fact that the semantic meanings of these codes are implicit, which poses challenges for interpretable, arbitrary, and flexible manipulation of expression by non-professionals. In this work, we show that it’s possible to convincingly alter a person’s expression in a user-friendly manner by offering localized control with adjustable intensity while preserving their identity and other attributes from the portrait. Our key insight is to employ action units (AUs) to represent facial expressions, and then steer a Stable-Diffusion model to produce high-quality expression editing results.
@@
In editing facial expressions, the crucial question is which representation should be used to encode expressions. Facial AUs, as anatomical markers of facial muscle activity, have proven effective for precise and flexible facial expression description in images [wang2019weakly, yan2022weakly, karaouglu2021self]. However, most AU annotations are of frontal faces of a limited number of subjects in laboratory settings, which easily to cause overfitting when used for training. A model trained by this type of data is hard to generalize to individuals with different poses and backgrounds, which we will present in Section IV-C. To deal with this issue, off-the-shelf AU intensity estimators like Libreface [chang2024libreface] used a knowledge distillation technique from a large-scale network pre-trained on natural images to accommodate the AU intensity estimation task on the lab datasets. Such automatic tools provide us with AU intensity estimation of high accuracy, hence we can produce AU estimation of any face image and use the estimation as an AU condition to train our model.
@@
On the other hand, diffusion models have emerged as a popular choice for image generation, surpassing Generative Adversarial Networks (GANs) with higher generation quality [zhang2023transformer, zhang2020supervised, xie2023consistency, yan2021fine]. ControlNet-style models enable users to add additional control signals, such as depth, paintings and skeleton pose, as the condition of the target [zhang2023adding]. Some works modify it to enable some basic expressions like happiness and sad [preechakul2021diffusion, zhang2023adding], but many of their editing results are with extreme face deformations that may look unrealistic. More importantly, these models can not provide specific controls over expressions, whether through text or through images. Such specific controls, including intensity and location of expressions, are the focus of our work.
@@
To merge the advantages of both worlds, we propose MagicFace, a model that allows us to correlate AU changes to facial expressions and then produce a photorealistic edited image conditioned on AU variations. Specifically, MagicFace first extracts AU intensities from portrait photographs using an off-the-shelf method, performs desired AU changes, and finally uses a Stable-Diffusion model to map the AU variations into photorealistic images. As the edited images should preserve the background and pose, we first cut out the background of the portrait and draw the contour of the pose, which is learned by an Attribute Controller, so that MagicFace only needs to perform conditional inpainting to the face. To maintain the consistency of identity and high-frequency facial characteristics, we introduce an ID encoder, which is a symmetrical UNet structure to capture spatial details of the input identity. This design allows the model to understand the relationship with the identity image within a uniform feature space, greatly enhancing the preservation of detailed appearance features. Our model is trained on a designed dataset with 30K image pairs. Fig. 1 shows the editing results of various people, background, and AU combinations.
In summary, our contributions are:
@@
  • An ID encoder capable of well preserving the attributes of the person in the edited portraits, unconstrained by any head poses, backgrounds, and characters.
  • Quantitative and qualitative results show that our method presents a more interpretable, flexible, and user-friendly editing manner with higher image generation quality compared to other facial expression editing methods.
@@ II Related Works
Figure 2: A display showcasing various action units and their corresponding intensity scales. Only a set of commonly used AUs is displayed here. For a complete collection of AUs with descriptions, see [ozelfaces].
@@ II-A Facial Action Units
The Facial Action Coding System (FACS) [ekman1978facial] is one of the most impactful methods for analyzing facial behavior. It is a comprehensive, anatomy-based system capable of encoding diverse facial movements through combinations of fundamental Action Units (AUs). AUs represent specific facial configurations resulting from the contraction of one or more facial muscles and are not influenced by emotional interpretation. The earlier version of FACS [ekman1978facial] included 44 Action Units (AUs), with 30 of them anatomically linked to specific facial muscles, while the remaining 14 were categorized as miscellaneous actions. In a later version [ekman2002facial], the criteria were updated: AU25, AU26, and AU27 were merged based on intensity, as were AU41, AU42, and AU43. The intensities of AUs can be assessed using a five-point ordinal scale, typically labeled with uppercase letters A to E, representing minimal to maximal intensity alongside the AU number. For example, AU2A indicates the minimum intensity of AU2, while AU2E represents its maximum intensity. However, this five-point scale is not uniform; for instance, levels C and D encompass a broader range of facial changes compared to levels A, B, and E. Fig. 2 shows example images labeled with specific AUs. The precision and clarity of AUs provide unmatched control over facial expressions, offering users an intuitive and easily interpretable set of tools, so we choose to use it as the condition to describe and edit facial expressions.
@@ II-B Facial Expression Editing
Facial expression editing is a challenging task as it demands a deep understanding of input facial images and prior knowledge about human expressions. Unlike general facial attribute editing, which primarily focuses on modifying the appearance of specific facial regions, facial expression editing is more complex as it often involves significant geometric transformations and simultaneous changes across multiple facial components. Remarkable progress has been made in earlier years with the advancement of GANs. ExprGAN [ding2018exprgan] introduces an expression controller to adjust the intensity of generated expressions, but it relies on a pre-trained face recognizer to maintain identity information. StarGAN [choi2018stargan], on the other hand, enables image translation across domains using a single model and preserves identity features by minimizing a cycle loss. However, it is limited to generating only discrete expressions. GANimation [pumarola2018ganimation] provides a more fine-grained method for editing facial expressions by leveraging Action Units (AUs). Similarly, ICface [tripathy2020icface] utilizes AUs to enable controllable facial expressions in facial reenactment. Recent focus has shifted to diffusion models due to their ability to produce high-quality images. However, many text-to-image diffusion models are unable to support flexible facial expression editing and, in most cases, can only handle simple emotion labels. Instead, some approaches choose the expression coefficients of 3DMM as the condition to enable fine-grained control over facial expression. However, the large number of coefficients poses a significant challenge, as only experts can manually adjust the desired expressions, which limits its applicability for general user-level applications. In contrast, this work focuses on injecting AUs into pre-trained large diffusion models to enable more flexible and user-friendly editing of the face.
@@ III Method
Figure 3: Overview of MagicFace. During training, a pair of images with the same identity but different pose, backgrounds, and expressions is used, respectively as the identity image and the target. AU variations are computed by an estimator and then sent into the denoising UNet as an AU condition. Pose and background of the target are parsed into an image condition independently, dealt with an Attribute Controller, and then input to the denoising UNet. ID encoder takes in the encoded identity image to edit for target AUs, where features in each transformer block are merged into the corresponding ones of the denoising UNet via self-attention. During inference, the conditional image will be parsed from the identity image.
@@ III-A Preliminariy
Stable Diffusion. Both the denoising UNet and ID encoder of our method inherit the same architecture as well as parameters of the denoising UNet in Stable Diffusion (SD). SD is developed from the latent diffusion model (LDM) [rombach2022high], which introduces learning of feature distributions to reduce computational complexity. It contains a fixed autoencoder to map input images into feature maps with a smaller size. Formally, for an image $\mathbf{x}$ given, the encoder maps it into a latent representation via $\mathbf{z}=\mathcal{E}(\mathbf{x})$ and the decoder reconstructs it by $\hat{\mathbf{x}}=\mathcal{D}(\mathbf{z})$. The denoising UNet in SD learns to denoise noise $\boldsymbol{\epsilon}$, which is normally distributed into $\mathbf{z}$. During training, the latent map $\mathbf{z}$ is added with Gaussian noise in $t$ timesteps and produces a noised latent $\mathbf{z}_{t}$. The UNet is trained to predict the noise added, optimized by the following objective:
@@
where $\epsilon$ denotes the denoising UNet and $c$ represents condition embeddings. In the original SD, $c$ is produced by a text encoder CLIP ViT-L/14 [alexey2020image] to enable text-to-image generations. During inference, $\mathbf{z}_{T}$ is sampled from a normal distribution as the noised latent at timestep $T$, and then it is denoised iteratively into $\mathbf{z}_{0}$ by a deterministic sampling process like DDPM [ho2020denoising], DDIM [song2020denoising]. Then $\mathbf{z}_{0}$ will be reconstructed by the decoder into a generated image.
AU variations VS absolute AUs. Stepping back and reconsidering the conditioning inputs, using absolute AUs should be an intuitive choice. However, this condition configuration has a drawback: the model must estimate the actual AUs of the input image to decide whether to edit it. From an application standpoint, we are required to supply a value that must exactly match the corresponding AU in the source image, even if no changes are intended. AU variations, as opposed to absolute AUs, represent the intended change in specific action units. This aligns with the definition of action units, which indicates the activation state of facial muscles. Therefore, in our model, we choose the difference between target AUs $\mathbf{c}_{tgt}$ and source AUs $\mathbf{c}_{ID}$ as the input condition, denoted as:
@@ III-B Architecture
Attribute Controller. As shown in Fig. 3, Attribute Controller takes in the latent representation of an image where the face is masked and the face pose is outlined by landmarks. This design is to solve the ill-posed problem: in videos shot in wild settings, there are no image pairs with identical backgrounds and poses, while expressions undergo significant changes for the same person, which means no ground truth of the image with target AUs is available. However, we can find numerous images of the same person with different backgounds, pose, and facial expressions. As an alternative, we separate the background and pose as an independent condition of the identity by parsing out the background image, which will prevent the model from learning all the information through the input identity image. On the parsed background image, we draw the contour to highlight the head pose in case some backgrounds may be black. It is theoretically possible that the denoising UNet will learn background and pose from the identity. Empirically, we find the network learns to use the Attribute Controller for background and pose information and does not rely on the ID encoder. We hypothesize this is because the conditional image is pixel-aligned with the ground truth and thus more easily used by the model.
@@
Attribute Controller is a convolution layer with 4 $\times$ 4 kernels, 2 $\times$ 2 strides, and 4 channels to align the latent of the conditional image with the same resolution as the noise input. The processed conditional image is then appended to the noisy latent together as the input to the denoising UNet.
@@
ID encoder. We first evaluated the vanilla ControlNet [zhang2023adding] for identity preservation. As illustrated in Fig. 8, we observed that ControlNet struggles to preserve appearance consistency when generating human images coherent to the AU prompt, making it unsuitable for our editing task. Many approaches employ the CLIP image encoder [radford2021learning] to encode image conditions, but in our experiments, it doesn’t adequately tackle problems associated with maintaining consistency in details. CLIP is designed to align semantic, high-level features with text; as a result, it will lose many details in the generated images. On the other hand, recent studies [xu2024ootdiffusion, hu2024animate, cao2023masactrl] have highlighted the significant role of self-attention layers in diffusion models in maintaining the details of generated images. Building on these findings, we conducted an experiment on self-attention for identity control. In this setup, both the identity image and the noisy image are processed through the self-attention layers. A key observation from this architecture is that it inherently facilitates appearance resemblance between the two images. One possible explanation is that the self-attention layers in the UNet play a crucial role in spatially transmitting appearance information. As a result, they can function as a deformation module, enabling the generation of visually similar images with varying geometric structures.
@@
Given the above observations, we design a feature extraction encoder $\omega_{\theta}$ for the identity image, which is aimed to preserve detailed face attributes, inspired by [hu2024animate]. This encoder shares the same architecture as the denoising UNet $\phi_{\theta}$. Specifically, let $\mathbf{s}=\mathcal{E}(\boldsymbol{I}_{id})$ denote the encoded latent feature sent to the ID encoder $\omega_{\theta}$ where $\boldsymbol{I}_{id}$ is the identity image. The features extracted from $\omega_{\theta}$ will be merged into the denoising UNet $\epsilon_{\theta}$ through self-attention. Along with AU conditions, the ID encoder and the denoising UNet are jointly optimized by:
@@
where $\mathbf{g}=\mathcal{E}(\boldsymbol{I}_{bg})$ denotes the encoded latent feature sent to the Attribute Controller $\phi_{\theta}$, $\boldsymbol{I}_{bg}$ is the background image and $\mathbf{c}_{AU}$ represents the encoded AU condition. We omit the AU encoder in the framework as it’s simply a linear layer to map the input AU into a feature with the same dimension as the time conditions in SD (details are deferred to Sec. IV-C). The AU condition is added to the time-embedding condition in $\phi_{\theta}$. To merge identity features from $\omega_{\theta}$ into $\phi_{\theta}$ with self-attention, we first dive into the transformer block [vaswani2017attention] of the ID encoder and denoising UNet, identifying each pair of feature maps that serve as inputs for the respective self-attention layers. For ease of explanation, we temporarily use $\mathbf{a}_{n}\in\mathbb{R}^{h_{n}\times w_{n}\times c_{n}}$ and $\mathbf{b}_{n}\in\mathbb{R}^{h_{n}\times w_{n}\times c_{n}}$ to denote the $n$th pair of feature maps from ID encoder and denoising UNet. In the denoising UNet, in each block the feature maps from both sides are first concatenated by $\mathbf{b}_{\mathbf{a}_{n}}=\mathbf{a}_{n}\bigoplus\mathbf{b}_{n}\in\mathbb{R}^{h_{n}\times 2w_{n}\times c_{n}}$, then we replace $\mathbf{b}_{n}$ with $\mathbf{b}_{\mathbf{a}_{n}}$ as the input to the self-attention layer. After coming out from the self-attention layer, we crop out the first half of the output feature map into $\mathbf{c}_{n}$ of resolution $h_{n}\times w_{n}\times c_{n}$ before input to the subsequent cross-attention layer.